On June 1, 2017, Royal Palm Company sold 6,000 of its 12%, 20-year, $1,000 face value bonds. At the time, the market rate for a similar bond is 8%. Interest payment dates are December 1 and June 1, and the company uses the effective rate method of amortization. On December 1, 2018, Royal Palm paid cash $4,250,000 to extinguish 4,000 of the bonds. At this time, the accrued interest was paid in cash. Prepare all necessary journal entries.
Answer #1

Journal:

Date Account Debit Credit
June 1, 2017 Cash $ 8,375,132.87
To Bonds payable $ 6,000,000.00
To premium on bonds payable $ 2,375,132.87
[Entry to record issue of bonds at premium]
#Refer note 1 for workings
Dec 1, 2017 Interest expense $     335,005.31
Premium on bonds payable $       24,994.69
To Cash $     360,000.00
[Entry to record interest expense, premium amortisation and interest payment]
#Refer note 2 for workings
Dec 31, 2017 Interest expense $       55,667.59
Premium on bonds payable $         4,332.41
To Interest payable $       60,000.00
[Entry to record interest accrual, amortisation of premium ]
#Refer note 3 for workings
June 1, 2018 Interest expense $     278,337.94
Interest payable $       60,000.00
Premium on bonds payable $       21,662.06
To Cash $     360,000.00
[Entry to record interest expense, premium amortisation and interest payment]
#Refer note 4 for workings
Dec 1, 2018 Interest expense $     332,965.75
Premium on bonds payable $       27,034.25
To Cash $     360,000.00
[Entry to record interest expense, premium amortisation and interest payment]
#Refer note 5 for workings
Dec 1, 2018 Bonds payable $ 4,000,000.00
Premium on bonds payable $ 1,531,406.30
To Cash $ 4,250,000.00
To Gain on extinguishment of debt $ 1,281,406.30
[ Entry to record extinguish 4000 bonds of total 6000, gain amount = bonds payable+premium - cash paid ]
#Refer note 6 for workings

Workings:

Note 1:

Bond price equals the discounted amounts of cash flows from the bond at the market rate of return.

Cash flows from the bonds are interest amounts and the maturity repayment amounts.

The rate used to discount these cash flows is 8% in this case, since interest is paid sem annually rate of interest for discount factors used is 8% * 6/12 = 4%.

Below is the computation to arrive at the price of the bonds.

Particulars Cash flow Discount factor Discounted cash flow
Interest payments-
Annuity factor (4%,40 periods)
360,000 19.7928               7,125,399
Principle payments - Present value factor (4%,40 periods) 6,000,000 0.2083               1,249,734
A Bond price         8,375,132.87
Face value               6,000,000
Premium/(Discount)         2,375,132.87
Interest amount:
Face value 6,000,000
Coupon/stated Rate of interest 12%
Frequency of payment(in months) 6
B Interest amount 6000000*0.12*6/12= 360,000
Present value calculation:
yield to maturity/Effective rate 8%
Effective interest per period(i) 0.08*6/12= 4%
Number of periods:
Ref Particulars Amount
a Number of interest payments in a year                                   2
b Years to maturiy                                 20
c=a*b Number of periods                                 40
# Note 2
Interest expense
Dec 1, 2017
Particulars Amount
Opening liability Incl premium = $                       8,375,132.87
Interest accrued for months = 6
Rate of interest required per annum = 8%
Interest expense = Liability * rate * months/12
= 8375132.87*0.08*6/12
= $                          335,005.31
Cash interest paid = $                          360,000.00
Amortization of bond premium:
Cash interest paid = $                          360,000.00
Less
Interest expense = $                         -335,005.31
Amortization of bond premium = $                             24,994.69
Closing liability amount:
Opening liability = $                       8,375,132.87
Less premium amortization = $                             24,994.69
Closing liability amount = $                       8,350,138.18
# Note 3
Interest expense
Dec 31, 2017
Particulars Amount
Opening liability Incl premium = $                       8,350,138.18
Interest accrued for months = 1
Rate of interest required per annum = 8%
Interest expense = Liability * rate * months/12
= 8350138.18*0.08*1/12
= $                             55,667.59
Cash interest accrued = $                             60,000.00
[6000000*0.12*1/12]
Amortization of bond premium:
Cash interest accrued = $                             60,000.00
Less
Interest expense = $                           -55,667.59
Amortization of bond premium = $                               4,332.41
# Note 4
Interest expense
June 1, 2018
Particulars Amount
Opening liability Incl premium = $                       8,350,138.18
Interest accrued for months = 5
Rate of interest required per annum = 8%
Interest expense = Liability * rate * months/12
= 8350138.18*0.08*5/12
= $                          278,337.94
Cash interest paid:
Cash interest accrued during the period = $                          300,000.00
[6000000*0.12*5/12]
Add: opening interest payable = $                             60,000.00
Total cash interest paid $                          360,000.00
Amortization of bond premium:
Cash interest accrued during the period = $                          300,000.00
Less
Interest expense = $                         -278,337.94
Amortization of bond premium = $                             21,662.06
Closing liability:
Opening liability = $                       8,350,138.18
Less
Premium amortisation till Dec.31 = $                               4,332.41
Premium amortisation till Jun.1 = $                             21,662.06
Closing liability = $                       8,324,143.71
# Note 5
Interest expense
Dec 1, 2018
Particulars Amount
Opening liability Incl premium = $                       8,324,143.71
Interest accrued for months = 6
Rate of interest required per annum = 8%
Interest expense = Liability * rate * months/12
= 8324143.71*0.08*6/12
= $                          332,965.75
Cash interest paid = $                          360,000.00
Amortization of bond premium:
Cash interest paid = $                          360,000.00
Less
Interest expense = $                         -332,965.75
Amortization of bond premium = $                             27,034.25
Closing liability amount:
Opening liability = $                       8,324,143.71
Less premium amortization = $                             27,034.25
Closing liability amount = $                       8,297,109.46
# Note 6
Apportionment of liability amounts:
Particulars Face value Premium Total liability
Of 6000 bonds $                             6,000,000 $ 2,297,109.46 $ 8,297,109.46
Per bond value $                                     1,000 $            382.85 $         1,382.85
Of 4000 bonds value $                             4,000,000 $ 1,531,406.30 $ 5,531,406.30

# Please comment, if any further explanation is required. Your rating is appreciated. #